# The Southern Photometrical Local Universe Survey (S-PLUS): searching for metal-poor dwarf galaxies

### v1

@@ Abstract
The metal content of a galaxy’s interstellar medium reflects the interplay between different evolutionary processes such as feedback from massive stars and the accretion of gas from the intergalactic medium. Despite the expected abundance of low-luminosity galaxies, the low-mass and low-metallicity regime remains relatively understudied. Since the properties of their interstellar medium resemble those of early galaxies, identifying such objects in the Local Universe is crucial to understand the early stages of galaxy evolution. We used the DR3 catalog of the Southern Photometric Local Universe Survey (S-PLUS) to select low-metallicity dwarf galaxy candidates based on color selection criteria typical of metal-poor, star-forming, low-mass systems. The final sample contains approximately 50 candidates. Spectral energy distribution fitting of the 12 S-PLUS bands reveals that 90% of the candidates are best fit by models with very low stellar metallicities. We obtained long-slit observations with the Gemini Multi-Object Spectrograph to follow-up a pilot sample and confirm whether these galaxies have low metallicities. We find oxygen abundances in the range 12 + log(O/H) (5% to 17% of the solar value), confirming their metal-poor nature. Most targets are outliers in the mass-metallicity relation, i.e. they display a low metal content relative to their observed stellar masses. In some cases, perturbed optical morphologies might give evidence of dwarf-dwarf interactions or mergers. These results suggest that the low oxygen abundances may be associated with an external event causing the accretion of metal-poor gas, which dilutes the oxygen abundance in these systems.
@@ 1 Introduction
Low-luminosity galaxies are less chemically evolved than more massive systems, likely due to a less efficient star formation activity and a higher fraction of metals lost in supernova events (Guseva et al., 2009; Kirby et al., 2013; McQuinn et al., 2015; Gonçalves, 2019). However, despite the expected large population of low-luminosity galaxies, the low-mass and low-metallicity regime is still relatively little studied among the galactic populations. In the standard cosmological scenario, dwarf galaxies play a crucial role in the hierarchical formation of structures, since the first objects to form and contribute, through mergers, to the growth of more massive systems were low-mass and metal-poor galaxies (Bovill & Ricotti, 2009; Bromm & Yoshida, 2011). Atek et al. (2024) have recently reported the discovery of eight faint galaxies in the redshift range between 6 and 8 with the James Webb Space Telescope. These galaxies exhibit stellar masses between and M⊙ and have oxygen abundances of 12 + log(O/H) 7, approximately 2% the solar value111In this work we adopt a solar oxygen abundance of 12 + log(O/H)⊙ = 8.69 0.04 (Asplund et al., 2021).. Nearby gas-rich, low-metallicity dwarfs are expected to be analogous to such primitive galaxies both because of their low abundances of heavy elements and the presence of intense bursts of star formation (Bouwens et al., 2016). However, due to their low intrinsic luminosity, detailed investigations of these objects at great distances is extremely complex. Therefore, it is important to combine the most detailed observations of the closest analogous objects with those of younger and more distant galaxies.
@@
Progress in identifying new metal-poor systems has been relatively slow. Detecting these faint galaxies requires that they are relatively nearby, and/or that they are experiencing a recent star formation episode. Robust abundance determinations are challenging to derive for faint galaxies, and because these systems are harder to find due to their low masses, the current metallicity relations are substantially underpopulated at the low-luminosity/low-mass end. Recent studies targeting large areas of the sky only managed to detect a small number of metal-poor dwarf galaxies with metallicity below 1/10th solar in the Local Universe (; Izotov et al., 2012; Hsyu et al., 2018; McQuinn et al., 2020; Lin et al., 2023). A few extremely metal poor (XMP) objects have been found at low redshifts with oxygen abundances around or slightly below 12 + log(O/H) (Izotov et al., 2009; Hirschauer et al., 2016; Izotov et al., 2018, 2019; Kojima et al., 2020; Izotov et al., 2024). These galaxies are characterized by having low stellar mass M M⊙ and very compact structures. Discovering and studying XMP galaxies is fundamental to understand the process of star formation in a metal poor ISM and to provide constraints for theoretical models of chemical evolution of low-mass galaxies.
@@
ID | S-PLUS ID | RA | DEC | zspec | log(M∗) | ||
[AB mag] | [AB mag] | [M⊙] | |||||
1 | DR4_3_SPLUS-s27s07_0011404 | 00:37:13.1 | -35:08:48 | 16.790.01 | 16.79 0.03 | 0.0138 | 8.500.11 |
2 | DR4_3_SPLUS-n14s01_0028263 | 10:00:27.6 | -17:28:25 | 17.540.02 | 17.51 0.04 | 0.0162 | 8.400.11 |
3 | DR4_3_SPLUS-n13s01_0024584 | 10:01:16.3 | -16:13:49 | 17.760.03 | 17.80 0.06 | 0.0075 | 7.530.11 |
4 | DR4_3_SPLUS-n02n27_0037849a | 12:23:23.3 | 01:48:52 | 17.070.02 | 17.00 0.04 | 0.0067 | 7.880.11 |
5 | DR4_3_SPLUS-s29s37_0046039 | 04:05:20.4 | -36:48:59 | 16.290.01 | 16.26 0.02 | 0.0027b | 7.340.11 |
6 | DR4_3_SPLUS-n03s22_0021537c | 11:57:12.0 | -02:41:12 | 17.100.02 | 17.20 0.04 | 0.0046d | 7.270.11 |
The Southern Photometric Local Universe Survey222https://splus.cloud/ (S-PLUS) is a 12-band optical survey conducted using the T80-South 0.83-m robotic telescope (T80S) located at Cerro Tololo Interamerican Observatory (Mendes de Oliveira et al., 2019). S-PLUS uses the Javalambre 12-band magnitude system (Cenarro et al., 2019), which includes four SDSS-like broad-band, a Javalambre intermediate-band filter, and seven narrow-band filters centered on prominent stellar spectral features such as [O ii]3727,3729, Caii H+K, H, H, and the Caii triplet. The panoramic camera features a single charge-coupled device (CCD) with 9.2k 9.2k pixels, a field of view of , and a pixel scale of 055 pix-1 (Benitez et al., 2014). The survey will cover an area of 9300 deg2 of the southern sky. Although there are other ongoing wide-area photometric surveys in the southern hemisphere, the S-PLUS unique 12 multi-band filter system enables a better sampling of the wavelength range between 3000 and 10000 to perform detailed analysis of the spectral energy distribution and stellar population of galaxies. Specifically, the availability of the and J0660333This narrow-band filter is centered at 6614 and is suitable for investigating the H emission line. filter imaging over a large-area of the sky makes S-PLUS a unique survey for studying star forming galaxies.
@@ 2 Sample Selection
Following H18, we applied color selection criteria typical of low-metallicity star-forming dwarf galaxies, such as IZw18 or Leo P. These color cuts were defined from the analysis of galaxies in the SDSS survey to identify low redshift ( 0.01–0.02), blue compact sources with colors consistent with H ii regions:
where is the uncertainty of the band magnitude, and is the filter covering the H+[N ii] lines. The fourth condition is defined to include also objects with a poorly constrained -band magnitude (H18). Taking advantage of the S-PLUS multiband filter set, we add an additional constraint, , to ensure that galaxies exhibit detectable H emission444We note that, given the characteristics of the narrow-band J0660 filter, the H emission can only be detected within the redshift range (Logroño-García et al., 2019; Mendes de Oliveira et al., 2019). (Gutiérrez-Soto et al., 2020). We applied these criteria to the S-PLUS DR3 catalog including only objects with mag ( 200000 objects). This selection returned a total of 380 candidate objects. We visually inspected the S-PLUS images to reject spurious candidates such as stars or star-forming regions located in the spiral arms of larger galaxies. After applying an absolute magnitude cut of mag555To derive the absolute magnitudes at this stage of the selection, we used the photometric redshifts derived by Lima et al. (2022) for the S-PLUS survey., which defines dwarf galaxies (Thuan & Martin, 1981; Meyer et al., 2014; Venhola et al., 2019), we obtained a final sample of 47 metal-poor dwarf galaxy candidates (see Fig. 1 as an example).
@@ 2.2 Spectral energy distribution (SED) fitting
CIGALE allows for a number of star formation histories (SFH) such as exponentially declined or delayed. We used a delayed exponential star formation history with e-folding times varying between and 7 Gyr for the older stellar populations and an exponential burst with e-folding time Myr for the latest star formation event. We left the stellar metallicity as a free parameter, ranging between 1/200th Z⊙ and 1 Z⊙, adopting the Calzetti et al. (2000) attenuation law. The majority of the candidates (43 out of 47) have SEDs that are best-fitted by stellar metallicity of Z = 0.0004 (1/50th Solar) or Z = 0.0001 (1/200th Solar). The stellar masses obtained from population synthesis models range between 107 and few times 109 M⊙, and they are comparable to those obtained using the relation between color and mass-to-light ratio derived by stellar population synthesis models (Bell et al., 2003)
where is the galaxy color corrected for the Milky Way extinction, absolute magnitude in AB mag units, also corrected for Milky Way extinction, and a Kroupa (2001) initial mass function (IMF) is assumed (see Table 1). This calibration has been used to derive the stellar masses of other samples of low-mass galaxies (e.g. H18, ). Therefore, we will use it throughout this work to facilitate comparison with other sets of low-metallicity dwarf galaxies (see Sections 2.3, 4.5, and 4.6). Following Bell et al. (2003) we adopt a systematic uncertainty of 25% for the mass-to-light ratios obtained with this method.
@@ 3 Gemini-S/GMOS observations and data reduction
The observations were carried out via two Fast Turnaround runs in 2022 and 2023 with the GMOS instrument (Hook et al., 2004) in the long slit mode on the Gemini-S telescope (Program IDs GS-2022B-FT-210, GS-2023A-FT-204). We used the B600 grism with a 2′′ slit width. The wavelength range varies between 4300 and 7500 Å with a central wavelength 5900 Å. This choice allowed us to include all the major emission lines needed to determine the oxygen abundance with the direct method, including the singly ionized ionic population of oxygen traced by the [O ii]7320,7330 lines. The spectral resolution estimated as the full-width-at-half-maximum (FWHM) of the emission lines of comparison lamps at 5900 is 9 , corresponding to = 195 km s-1 ( = FWHM/2.355). The total integration time ranged between 1350 and 7200 seconds. Data were binned by 2 in the spatial direction and binned by 2 in the spectral direction only for the first galaxy observed in the 2022 observing run (ID 1). The central position of the long slit was defined according to the peak of H emission in the S-PLUS continuum-subtracted J0660 images of each galaxy. The position angle of the slit for the first target (ID 1) was set at the parallactic angle. For the remaining targets, however, the slit was oriented to cover most of the galactic disc and the H ii regions detected in the J0660 narrow-band images (Fig. 3, top panels). Spectra of CuAr arc lamps were obtained for wavelength calibrations, along with bias frames and dome flats to correct for the detector bias level and pixel-to-pixel variations, respectively. The spectrophotometric standard LTT 2415 was observed with the same experimental set-up and it was used to calibrate the targets in relative flux.
@@ 4.1 Emission line fluxes and intensities
In Fig. 4 we display the 1D spectra obtained for the four objects, with the prominent emission lines marked. The brightest lines are H and H recombination lines, along with the [O iii]4959,5007 lines. As can be seen in the figure, even with long exposures on a 8-meter telescope, the detection of the auroral line is challenging in these galaxies. We measured the centers, the FWHM, the equivalent widths (EW), and the fluxes of the detected emission lines by performing Gaussian fitting of the line profiles with the PySpecKit (Ginsburg et al., 2022) python package.
The emission line fluxes were corrected for internal extinction using the Balmer decrement. The Balmer decrement remains roughly constant for typical gas conditions in star-forming galaxies (Osterbrock & Ferland, 2006), and we assumed the theoretical values for Case B recombination and 104 K, [(H)/(H)]th = 2.86 and [(H)/(H)]th = 0.47 (Osterbrock & Ferland, 2006). This choice is standard for star-forming galaxies in the literature. Internal reddening was derived from the comparison of the H/H and H/H ratios with their theoretical values. The increase of the Balmer decrement over the intrinsic value due to the extinction can be expressed in terms of the Balmer color excess (Momcheva et al., 2013):
where [(H)/(H)]obs is the observed ratio of the two fluxes. The extinction at the H wavelength is related to the Balmer color excess as following:
where is the assumed extinction curve and the extinction coefficient is defined as . For our targets, we used the Small Magellanic Cloud (SMC) extinction law (Prevot et al., 1984)777This extinction curve is the one that is implemented by default in the tool that we used to determine the ionized gas properties (see Sect. 4.3). Using this curve instead of the more recent extinction measurements by Gordon et al. (2003) does not significantly affect our results. which provides a good representation of a metal-poor interstellar medium.

### v2

@@ Abstract
The metal content of a galaxy’s interstellar medium reflects the interplay between different evolutionary processes such as feedback from massive stars and the accretion of gas from the intergalactic medium. Despite the expected abundance of low-luminosity galaxies, the low-mass and low-metallicity regime remains relatively understudied. Since the properties of their interstellar medium resemble those of early galaxies, identifying such objects in the Local Universe is crucial to understand the early stages of galaxy evolution. We used the DR3 catalog of the Southern Photometric Local Universe Survey (S-PLUS) to select low-metallicity dwarf galaxy candidates based on color selection criteria typical of metal-poor, star-forming, low-mass systems. The final sample contains approximately 50 candidates. Spectral energy distribution fitting of the 12 S-PLUS bands reveals that 60% of the candidates are best fit by models with low stellar metallicities. We obtained long-slit observations with the Gemini Multi-Object Spectrograph to follow-up a pilot sample and confirm whether these galaxies have low metallicities. We find oxygen abundances in the range 12 + log(O/H) (4% to 13% of the solar value), confirming their metal-poor nature. Most targets are outliers in the mass-metallicity relation, i.e. they display a low metal content relative to their observed stellar masses. In some cases, perturbed optical morphologies might give evidence of dwarf-dwarf interactions or mergers. These results suggest that the low oxygen abundances may be associated with an external event causing the accretion of metal-poor gas, which dilutes the oxygen abundance in these systems.
@@ 1 Introduction
Low-luminosity galaxies are less chemically evolved than more massive systems, likely due to a less efficient star formation activity and a higher fraction of metals lost in supernova events (Guseva et al., 2009; Kirby et al., 2013; McQuinn et al., 2015; Gonçalves, 2019). However, despite the expected large population of low-luminosity galaxies, the low-mass and low-metallicity regime is still relatively little studied among the galactic populations. In the standard cosmological scenario, dwarf galaxies play a crucial role in the hierarchical formation of structures, since the first objects to form and contribute, through mergers, to the growth of more massive systems were low-mass and metal-poor galaxies (Bovill & Ricotti, 2009; Bromm & Yoshida, 2011). Chemerynska et al. (2024) have recently reported the discovery of eight faint galaxies in the redshift range between 6 and 8 with the James Webb Space Telescope. These galaxies exhibit stellar masses between and M⊙ and have oxygen abundances of 12 + log(O/H) 7, approximately 2% the solar value111In this work we adopt a solar oxygen abundance of 12 + log(O/H)⊙ = 8.69 0.04 (Asplund et al., 2021).. Nearby gas-rich, low-metallicity dwarfs are expected to be analogous to such primitive galaxies both because of their low abundances of heavy elements and the presence of intense bursts of star formation (Bouwens et al., 2016). However, due to their low intrinsic luminosity, detailed investigations of these objects at great distances is extremely complex. Therefore, it is important to combine the most detailed observations of the closest analogous objects with those of younger and more distant galaxies.
@@
Progress in identifying new metal-poor systems has been relatively slow. Detecting these faint galaxies requires that they are relatively nearby, and/or that they are experiencing a recent star formation episode. Robust abundance determinations are challenging to derive for faint galaxies, and because these systems are harder to find due to their low masses, the current metallicity relations are substantially underpopulated at the low-luminosity/low-mass end. Recent wide-field surveys have detected only a small number of dwarf galaxies with metallicities below one tenth of solar (commonly referred to as metal-poor dwarfs, Kunth & Östlin (2000)) in the Local Universe (; Izotov et al., 2012; Hsyu et al., 2018; McQuinn et al., 2020; Lin et al., 2023). A few extremely metal poor (XMP) objects (i.e. with 12 + log(O/H) 7.35, approximately 5% of the solar value, McQuinn et al. (2020)) have been found at low redshifts (Izotov et al., 2009; Hirschauer et al., 2016; Izotov et al., 2018, 2019; Kojima et al., 2020; Izotov et al., 2024a). These galaxies are characterized by having low stellar mass M M⊙ and very compact structures. Discovering and studying XMP galaxies is fundamental to understand the process of star formation in a metal poor ISM and to provide constraints for theoretical models of chemical evolution of low-mass galaxies.
@@
| ID | S-PLUS ID | RA | DEC | zspec | log(M∗) | |||
| (J2000) | (J2000) | [AB mag] | [AB mag] | [AB mag] | [M⊙] | |||
| 1 | DR4_3_SPLUS-s27s07_0011404 | 00:37:13.1 | -35:08:48 | 16.99 0.02 | 16.790.01 | 16.79 0.03 | 0.0138 | 8.500.11 |
| 2 | DR4_3_SPLUS-n14s01_0028263 | 10:00:27.6 | -17:28:25 | 17.84 0.05 | 17.540.02 | 17.51 0.04 | 0.0162 | 8.400.11 |
| 3 | DR4_3_SPLUS-n13s01_0024584 | 10:01:16.3 | -16:13:49 | 18.14 0.05 | 17.760.03 | 17.80 0.06 | 0.0075 | 7.530.11 |
| 4 | DR4_3_SPLUS-n02n27_0037849a | 12:23:23.3 | 01:48:52 | 17.18 0.03 | 17.070.02 | 17.00 0.04 | 0.0067 | 7.880.11 |
| 5 | DR4_3_SPLUS-s29s37_0046039 | 04:05:20.4 | -36:48:59 | 16.39 0.01 | 16.290.01 | 16.26 0.02 | 0.0027b | 7.340.11 |
| 6 | DR4_3_SPLUS-n03s22_0021537c | 11:57:12.0 | -02:41:12 | 17.28 0.03 | 17.100.02 | 17.20 0.04 | 0.0046d | 7.270.11 |
The Southern Photometric Local Universe Survey222https://splus.cloud/ (S-PLUS) is a 12-band optical survey conducted using the T80-South 0.83-m robotic telescope (T80S) located at Cerro Tololo Interamerican Observatory (Mendes de Oliveira et al., 2019). S-PLUS uses the Javalambre 12-band magnitude system (Cenarro et al., 2019), which includes four SDSS-like broad-band, a Javalambre intermediate-band filter, and seven narrow-band filters centered on prominent stellar spectral features such as [O ii]3727,3729, Caii H+K, H, H, and the Caii triplet. The panoramic camera features a single charge-coupled device (CCD) with 9.2k 9.2k pixels, a field of view of , and a pixel scale of 055 pix-1 (Benitez et al., 2014). The survey will cover an area of 9300 deg2 of the southern sky. Although there are other ongoing wide-area photometric surveys in the southern hemisphere, the S-PLUS unique 12 multi-band filter system enables a better sampling of the wavelength range between 3000 and 10000 to perform detailed analysis of the spectral energy distribution and stellar population of galaxies. Specifically, the availability of the and J0660333This narrow-band filter is centered at 6614 and is suitable for investigating the H emission line. filter imaging over a large-area of the sky makes S-PLUS a unique survey for studying star forming galaxies. Moreover, it has been shown that an improved accuracy in physical parameter estimation from spectral energy distribution fitting (SED) can be attained when combining medium- and narrow-band filters with broad-band ones (Díaz-García et al., 2015; Mejía-Narváez et al., 2017; San Roman et al., 2019; Molino et al., 2020; González Delgado et al., 2021; Thainá-Batista et al., 2023; Nakazono et al., 2024).
@@ 2 Sample Selection
Following H18, we applied color selection criteria typical of low-metallicity star-forming dwarf galaxies, such as IZw18 or Leo P. These color cuts were adapted from the analysis of galaxies in the SDSS survey to identify low redshift ( 0.01–0.02), blue compact sources with colors consistent with H ii regions:
where is the uncertainty of the band magnitude, and is the filter covering the H+[N ii] lines. The fourth condition is defined to include also objects with a poorly constrained -band magnitude (H18). Taking advantage of the S-PLUS multiband filter set, we add an additional constraint, , to ensure that galaxies exhibit detectable H emission444We note that, given the characteristics of the narrow-band J0660 filter, the H emission can only be detected within the redshift range (Logroño-García et al., 2019; Mendes de Oliveira et al., 2019). (Gutiérrez-Soto et al., 2020). Figure 1 shows the SDSS spectrum of one of the galaxies of our sample, with the S-PLUS photometry and transmission curves of the 12 filter system overlaid. We applied these criteria to the S-PLUS DR3 catalog including only objects with mag ( 200000 objects). This selection returned a total of 380 candidate objects. We visually inspected the S-PLUS images to reject spurious candidates such as stars or star-forming regions located in the spiral arms of larger galaxies. After applying an absolute magnitude cut of mag555To derive the absolute magnitudes at this stage of the selection, we used the photometric redshifts derived by Lima et al. (2022) for the S-PLUS survey., which defines dwarf galaxies (Thuan & Martin, 1981; Meyer et al., 2014; Venhola et al., 2019), we obtained a final sample of 47 metal-poor dwarf galaxy candidates (see Fig. 2 as an example).
@@ 2.2 Spectral energy distribution (SED) fitting
CIGALE allows for a number of star formation histories (SFH) such as exponentially declined or delayed. We used a delayed exponential star formation history with e-folding times varying between and 7 Gyr for the older stellar populations and an exponential burst with an age of 20 Myr and e-folding time Myr for the latest star formation event. We left the stellar metallicity as a free parameter, ranging between 1/200th Z⊙ and 1 Z⊙, adopting the Small Magellanic Cloud (SMC) attenuation law (Boquien et al., 2019). The majority of the candidates (29 out of 47) have SEDs that are best-fitted by stellar metallicity below Z = 0.002 (1/10th Solar). The stellar masses obtained from population synthesis models range between 107 and few times 109 M⊙, and they are comparable to those obtained using the relation between color and mass-to-light ratio derived by stellar population synthesis models (Bell et al., 2003)
where is the galaxy color corrected for the Milky Way extinction, absolute magnitude in AB mag units, also corrected for Milky Way extinction, and a Kroupa (2001) initial mass function (IMF) is assumed (see Table 1). This calibration has been used to derive the stellar masses of other samples of low-mass galaxies (e.g. H18, ). Therefore, we will use it throughout this work to facilitate comparison with other sets of low-metallicity dwarf galaxies (see Sections 2.3, 4.7, and 4.8). Following Bell et al. (2003) we adopt a systematic uncertainty of 25% for the mass-to-light ratios obtained with this method.
@@ 3 Gemini-S/GMOS observations and data reduction
The observations were carried out via two Fast Turnaround runs in 2022 and 2023 with the GMOS instrument (Hook et al., 2004) in the long slit mode on the Gemini-S telescope (Program IDs GS-2022B-FT-210, GS-2023A-FT-204). We used the B600 grism with a 2′′ slit width. The wavelength range varies between 4300 and 7500 Å with a central wavelength 5900 Å. This choice allowed us to include all the major emission lines needed to determine the oxygen abundance with the direct method, including the singly ionized ionic population of oxygen traced by the [O ii]7320,7330 lines. The spectral resolution estimated as the full-width-at-half-maximum (FWHM) of the emission lines of comparison lamps at 5900 is 9 , corresponding to = 195 km s-1 ( = FWHM/2.355). The total integration time ranged between 1350 and 7200 seconds. Data were binned by 2 in the spatial direction and binned by 2 in the spectral direction only for the first galaxy observed in the 2022 observing run (ID 1). The central position of the long slit was defined according to the peak of H emission in the S-PLUS continuum-subtracted J0660 images of each galaxy. The position angle of the slit for the first target (ID 1) was set at the parallactic angle. For the remaining targets, however, the slit was oriented to cover most of the galactic disk and the H ii regions detected in the J0660 narrow-band images (Fig. 4, top and central panels). With a slit width of 2′′ generously exceeding the characteristic seeing, and all observations conducted under favorable low-airmass conditions (average value 1.4), our setup was inherently robust against the minimal effects of atmospheric differential refraction, guaranteeing the photometric integrity of the resulting spectra (Filippenko, 1982). Spectra of CuAr arc lamps were obtained for wavelength calibrations, along with bias frames and dome flats to correct for the detector bias level and pixel-to-pixel variations, respectively. The spectrophotometric standard LTT 2415 was observed with the same experimental set-up and it was used to calibrate the targets in relative flux.
@@ 4.1 Emission line fluxes and intensities
In Fig. 5 we display the 1D spectra obtained for the four objects, with the prominent emission lines marked. The brightest lines are H and H recombination lines, along with the [O iii]4959,5007 lines. As can be seen in the figure, even with long exposures on a 8-meter telescope, the detection of the auroral line is challenging in these galaxies.
4.2 Continuum subtraction and line extraction with ppXF.
We used the penalized pixel-Fitting (pPXF) (Cappellari & Emsellem, 2004; Cappellari, 2017, ppXF, ) algorithm to model the stellar continuum and extract the emission line fluxes. pPXF was applied to the four GMOS target and to galaxy No. 6 which has a SDSS spectrum. The stellar continuum was modeled to account for absorption features present in some of the targets, reducing their impact on the measurement of the emission line fluxes (e.g. Balmer absorption). The fit was performed using single stellar population (SSP) templates from the E-MILES library (Vazdekis et al., 2016; Cappellari, 2023). pPXF constructs an optimal combination of these templates and fits them to the observed spectrum using Gauss–Hermite expansions for the line-of-sight velocity distribution. Multiplicative polynomials of degree 8–10 (depending on the galaxy) were employed to match the overall spectral shape of the data. The emission lines were simultaneously fitted together with the continuum and the line-of-sight velocity was derived by fitting the emission lines simultaneously, assuming a fixed velocity width. The instrumental resolution was taken into account when setting the initial velocity dispersion. The uncertainties on the emission-line fluxes were taken directly from the formal errors provided by pPXF, which are derived from the covariance matrix of the fit and account for both the noise in the spectrum and the coupling between fitted parameters. The complete list of the emission lines that were fit in each galaxy is given in Table 2.
4.3 Extinction correction
The measured emission-line fluxes were first corrected for galactic extinction using the Cardelli et al. (1989) law and then corrected for internal extinction with the Nebular Empirical Analysis Tool (NEAT, Wesson et al., 2012), determined from the ratios of hydrogen Balmer lines in an iterative procedure. The extinction coefficient H() is initially estimated using the observed Balmer line ratios (H/H and H/H) compared to their theoretical values at the electron temperature = 10000 K and density = 1000 cm-3. A flux-weighted average of the individual ratios is adopted as the first estimate of H(), which is then used to deredden the line fluxes. Electron temperatures and densities are derived from the dereddened spectrum (as described in Sect. 4.5), and a new value of H() is recalculated using the corresponding theoretical Balmer ratios. This procedure is iterated until H() converges (Wesson et al., 2012). The extinction coefficient for our sample was derived using the Small Magellanic Cloud (SMC) extinction law, (Prevot et al., 1984)777This extinction curve is the one that is implemented by default in the tool that we used to determine the ionized gas properties (see Sect. 4.5). Using this curve instead of the more recent extinction measurements by Gordon et al. (2003) does not significantly affect our results, which provides a good representation of a metal-poor interstellar medium. In most of our targets, the observed Balmer decrement (H/H) was lower than the theoretical value expected for case B recombination (Atek et al., 2009; Yang et al., 2017; Scarlata et al., 2024). With the exception of target No. 3, the extinction coefficients derived from the H/H ratios are small suggesting the presence of low internal extinction (Table 3).